A current of 90 A is shared by three resistances in parallel. The wires are of the


same material and have their lengths in the ratio 2: 3: 4 and their cross-sectional areas


in the ratio 1: 2: 3. Determine current in each resistance

1
Expert's answer
2022-01-17T08:00:04-0500

The voltage across all resistances is the same:


"V=I_1R_1=I_2R_2=I_3R_3."


By Kirchhoff's current law:


"I_1+I_2+I_3=90\text{ A}.\\\space\\
\dfrac V{R_1}+\dfrac V{R_2}+\dfrac V{R_3}=90\text{ A}."


The resistance of each resistor is


"R_1=\dfrac{\rho L_1}{\pi r_1^2},\\\space\\
R_2=\dfrac{\rho L_2}{\pi r_2^2}=\dfrac{\rho (\frac 32L_1)}{\pi (2r_1)^2}=\dfrac 38 R_1,\\\space\\
R_3=\dfrac{\rho L_3}{\pi r_3^2}=\dfrac{\rho (2L_1)}{\pi (3r_1)^2}=\dfrac 29 R_1."


Since "I_1R_1=I_2R_2=I_3R_3,"


"I_1=\dfrac 38I_2=\dfrac29I_3,\\\space\\
I_1+\dfrac 83 I_1+\dfrac 92I_1=90,\\\space\\
I_1=11\text{ A},\\
I_2=29\text{ A},\\
I_3=50\text{ A}."